### v2

# Tredoku Patterns

Abstract
Donald A. Preece gave two talks in 2013, in which he introduced the notion of tredoku patterns. These are certain configurations of diamond-shaped tiles, inspired by the sukodu-like puzzle that appears in newspapers. Very sadly, Prof.  Preece died in January 2014, before he was able to complete his work on this problem. This paper reconstructs his work, and settles (in the affirmative) the conjectures he proposed.
1 Introduction
As the reader knows, sudoku is a game based on a $3\times 3$ grid of squares. Each square is subdivided into nine smaller square entries, some of which are filled with the numbers 1 to 9. The aim is to fill in the blank entries so that the entries in any square, any row (crossing a horizontal run of three squares) and any column (crossing a vertical run of three squares) form a permutation of $\{1,2,\ldots,9\}$. Tredoku [^1] Tredoku is a trademark of Mindome Ltd. Tredoku puzzles regularly appear in The Times (of London), for example. is a variant of sudoku invented by Eyal Amitzur, where the squares are replaced by a collection of diamonds (see, for example, Figure 1); rows might change direction as they cross two adjacent diamonds that share an edge. Tredoku is presented as a 3-dimensional version of sudoku, with the diamonds shaded so that the region appears to be part of a surface of an object made from cubes [6]. Latex for the display of Tredoku puzzles has been explored in [9].
Donald Preece asked: What arrangements of diamonds might be used in tredoku puzzles? He called these arrangements tredoku patterns. The question can be thought of as being part of the theory of diamond tilings (often called lozenge tilings) studied as part of combinatorial geometry, algebraic combinatorics, and physics. The history of such tilings in combinatorics goes back as least as far as 1896, where Percy MacMahon [4] conjectured a formula for the number of diamond tilings of a hexagonal region of the plane; he established this formula in 1916 [5]. In physics, diamond tilings are studied as simple instantiations of the 3-dimensional Ising model; in particular there have been some beautiful results on random tilings motivated by this model [2]. Diamond tilings, and the related domino tilings, are the basic examples which are generalised by the theory of dimers [3]. One interesting aspect of tredoku patterns is the introduction of the concept of a run of tiles (which we define below) to this research area.
Donald Preece gave two talks [7, 8] on this topic in 2013 in which he defined tredoku patterns, and provided many examples. (The precise definition is below. In Appendix A, the abstract from [7] is reproduced, which contains the definition in his own words.) In these talks, he provided various constructions for tredoku patterns, and proposed questions and conjectures concerning which parameters are possible. Prof Preece died in January 2014, before he was able to complete his work on tredoku patterns. This paper solves (in the affirmative) the conjectures he proposed.
I am working from the notes I took when attending his talk in July 2013 [7]. Prof.  Preece distributed copies of his notes on the topic after his talk in October [8], in order to encourage attendees to finish his work. Unfortunately, I did not take a copy myself, and I am unable to locate any remaining copies. If anyone reading this paper does have a copy of these notes, please get in touch with me: I would be very interested to see their contents.
In the remainder of the introduction, I will define tredoku patterns, and the parameters we are considering. I then state the main theorem of the paper, which characterises the parameters where tredoku patterns exist, positively resolving Preece’s conjectures.
Definition.
Consider a tiling of the plane using equilateral triangles, meeting along their edges. A pattern $\mathcal{T}$ of diamond-shaped tiles is a finite set of non-overlapping tiles, each the union of two triangles from our tiling that meet along an edge.
Figure 1 is an example of a pattern. We see that each tile can be oriented in one of three ways (shaded differently in the figure).
Definition.
A run of length $\ell$ in a pattern $\mathcal{T}$ is a maximal sequence $t_{1},t_{2},\ldots,t_{\ell}$ of distinct tiles in $\mathcal{T}$ such that adjacent tiles in the sequence are edge-adjacent, and these edges are all parallel.
Figure 1: A tredoku pattern
Definition.
A tredoku pattern $\mathcal{T}$ is a pattern of $3$ or more diamond-shaped tiles that satisfies the following four conditions:
  1. 1.
    If there exists a run of length $\ell$ in $\mathcal{T}$, then $\ell\in\{1,3\}$.
  2. 2.
    The pattern is edge-connected. More precisely: Define a graph $\Gamma(\mathcal{T})$ whose vertices correspond to the tiles in $\mathcal{T}$, with vertices being adjacent if and only if the correponding tiles share an edge. Then $\Gamma(\mathcal{T})$ is connected.
  3. 3.
    The pattern has no ‘holes’. More precisely: Define $|\mathcal{T}|$ to be the region of the plane covered by the tiles in $\mathcal{T}$. Then $|\mathcal{T}|$ is simply connected.
  4. 4.
    If any tile of the pattern $\mathcal{T}$ is removed, $\mathcal{T}$ remains connected (possibly via the corners of tiles touching at a vertex). More precisely: For any tile $t\in\mathcal{T}$, the region $|\mathcal{T}\setminus\{t\}|$ of the plane is path connected for every $t\in\mathcal{T}$.
So the pattern in Figure 1 is a tredoku pattern with $7$ tiles, $4$ runs of length $3$, and $2$ runs of length $1$.
Definition.
A leaf in a tredoku pattern is a tile that is contained in exactly one run of length $3$ (and one run of length $1$, since every tile is contained in two runs).
For example, the pattern in Figure 1 contains $2$ leaves. We have the following lemma:
Lemma 1.
Preece [7, 8]. If there exists a tredoku pattern with $\tau$ tiles, $\rho$ runs and $\ell$ leaves then $0\leq\ell\leq\tau$ and $\ell=2\tau-3\rho$.
Proof.
Since every leaf is a tile, it is clear that $0\leq\ell\leq\tau$. Counting, in two ways, pairs $(t,r)$ where $t$ is a tile in the pattern and $r$ is a run of length $3$, we find that $3\rho=2(\tau-\ell)+\ell$, and so $\ell=2\tau-3\rho$. ∎
Preece’s conjectures [7] were:
  • (a)
    There does not exist a $12$ tile tredoku pattern with 00 leaves (and $8$ runs of length $3$).
  • (b)
    There does not exist a $15$ tile tredoku pattern with $9$ leaves (and $7$ runs of length $3$).
  • (c)
    For $\rho\geq 9$, there does not exist a tredoku pattern with $\rho$ runs of length $3$, $12\rho+$ 1 tiles and $\rho+2$ leaves.
We see that all of these conjectures are corollaries of the following theorem, which characterises the parameters $\tau$, $\rho$ and $\ell$ such that a $\tau$-tile treduku pattern with $\rho$ runs and $\ell$ leaves exists.
Theorem 2.
Let $\tau$, $\rho$ and $\ell$ be non-negative integers with $\tau\geq 3$, such that $\ell\leq\tau$ and $\ell=2\tau-3\rho$.
  • (i)
    If $\ell>\lceil\tau/2\rceil+1$, then a $\tau$-tile tredoku pattern with $\rho$ runs of length $3$ and $\ell$ leaves does not exist.
  • (ii)
    Suppose
    $\begin{split}(\tau,\rho,\ell)\in\{(3,1,3),(3,2,0),(4,2,2),(5,3,1),(6,4,0),\\ (12,8,0),(15,7,9)\}\end{split}$
    or $(\tau,\rho,\ell)$ is of the form $(2\rho+1,\rho,\rho+2)$ with $\rho\geq 9$. Then a $\tau$-tile tredoku pattern with $\rho$ runs of length $3$ and $\ell$ leaves does not exist.
  • (iii)
    If neither of the conditions (i) and (ii) hold, then a $\tau$-tile tredoku pattern with $\rho$ runs of length $3$ and $\ell$ leaves exists.
Theorem 2 is visualised by Figure 2, where we plot $\tau$ against $\ell$, and where the shaded squares are parameters where tredoku patterns exist. I believe Donald Preece had a proof of Theorem 2 modulo his conjectures above, so the material in Sections 23 and 4 below is a reconstruction of his work. His conjectures are solved in Sections 5 and 6.
[…]
Figure 2: Parameters $(\tau,\ell)$ where tredoku patterns exist are drawn as shaded squares. We must have $\rho=(2\tau-\ell)/3$.
The remainder of the paper is organised as follows. Theorem 2 (iii) is proved in Section 2, by providing examples of tredoku patterns for all the possible parameters listed there. Section 3 proves Theorem 2 (i). Condition 4 of being a tredoku pattern is not always easy to check when only local information about the pattern is known; Section 4 provides an equivalent condition which is easier to check. Section 5 classifies all tredoku patterns with $\rho$ runs of length $3$, with $12\rho+$ 1 tiles and $\rho+2$ leaves. In particular, the results in Section 5 show that $\tau$-tile tredoku patterns with $\rho$ runs and $\ell$ leaves do not exist when $(\tau,\rho,\ell)\in\{(3,1,3),(15,7,9)\}$ or when $(\tau,\rho,\ell)=(2\rho+1,\rho,\rho+2)$ with $\rho\geq 9$. Section 6 completes the proof of Theorem 2, by showing that tredoku patterns do not exist in the cases when $(\tau,\rho,\ell)\in\{(2,1,1),(3,2,0),(4,2,2),(5,3,1),(6,4,0),(12,8,0)\}$. Most of these cases are not at all hard, but the case when $(\tau,\rho,\ell)=(12,8,0)$ requires a more detailed argument. Section 7 is a short conclusion to the paper.
This paper is dedicated to the memory of Donald Preece. He was a fund of interesting problems and anecdotes, and a pleasure to know. Rosemary Bailey has written a summary of his work, which can be found on arXiv [1], which I recommend very highly.
2 Examples of tredoku patterns
This section provides examples of tredoku patterns for many combinations of parameters; this establishes Theorem 2 (iii).
First, consider tredoku patterns with $\tau$ tiles and no leaves; these are the parameters in the bottom row of Figure 2. Theorem 2 asserts that such patterns occur when $\tau=9$, or when $\tau$ is a multiple of $3$ with $\tau\geq 15$. Examples for $\tau=9$ and $\tau=15$ are drawn in Figure 3, and examples for $\tau=18+3i$ for $i=0,1,2$ are drawn in Figure 4. Examples when $\tau=18+3i$ for any larger integer $i$ may be constructed by extending the central repeating section of the pattern.
Figure 3: Tredoku patterns with $9$ and $15$ tiles, and no leaves.
Figure 4: Tredoku patterns with $18+3i$ tiles and no leaves, for $i=0,1,2$.
The tredoku patterns drawn in Figures 5 and 6, together with patterns obtained from Figure 6 by extending the repeating central section, give examples of tredoku patterns for the one-leaf parameters claimed by Theorem 2. Similarly, Figures 7 and 8 give examples for all two-leaf parameters, Figures 9 and 10 give examples for all three-leaf parameters, and Figures 11 and 12 give examples for all four-leaf parameters. These are the parameters for the bottom five rows in Figure 2.
Figure 5: Tredoku patterns with $8$, $11$, $14$ and $17$ tiles, and one leaf.
Figure 6: Tredoku patterns with $20+3i$ tiles and one leaf, for $i=0,1,2$.
Figure 7: Tredoku patterns with $7$ and $10$ tiles, and two leaves.
Figure 8: Tredoku patterns with $13+3i$ tiles and two leaves, for $i=0,1,2$.
Figure 9: Tredoku patterns with $6$, $9$ and $12$ tiles, and three leaves.
Figure 10: Tredoku patterns with $15+3i$ tiles and three leaves, for $i=0,1,2$.
Figure 11: Tredoku patterns with $5$ tiles and $9$ tiles and four leaves. One leaf is shaded in the $9$ tile pattern.
Figure 12: Tredoku patterns with $11+3i$ tiles and four leaves, for $i=0,1,2$, with one leaf shaded.
Starting with one of our examples of a four-leaf tredoku pattern, provided our example has $\tau$ tiles with $\tau>5$, we may extend outwards from the shaded leaf by adding $2j$ tiles as described in Figure 13. The result is a pattern with $\tau+2j$ tiles and $4+j$ leaves. (This process does not work for $\tau=5$, as none of the leaves in our example have the correct from.) We have now covered all parameters required by Theorem 2 (iii), apart from the five sets of parameters with $\tau>5$ which lie on the upper diagonal of Figure 2. These five sets of parameters are covered by the examples in Figure 14. This completes the proof of Theorem 2 (iii).
Figure 13: Adding $2j$ tiles and $j$ runs to a four-leaf tredoku pattern, for $j=1,2,3,4.$ The four-leaf pattern contains the shaded leaf, and lies to the left of the dashed line.
Figure 14: Tredoku patterns with $\rho$ runs of length $3$, $12\rho+$ 1 tiles and $\rho+2$ leaves, for $\rho=3,4,5,6$ and $8$.
3 Tredoku patterns with many leaves
In this section, we aim to prove Theorem 2 (i). To this end, we make the following definition.
Definition.
A weak tredoku pattern is a pattern of $3$ or more tiles that satisfies Conditions 1 to 3 in the definition of a tredoku pattern, but does not necessarily satisfy Condition 4.
We note that the argument of Lemma 1 does not require Condition 4, so a weak tredoku pattern with $\tau$ tiles, $\rho$ runs of length $3$ and $\ell$ leaves must satisfy the equality $\ell=2\tau-3\rho$. It is easy to see that there are no weak tredoku patterns with $2$ tiles, and that a weak tredoku pattern with $3$ tiles consists of a single run of three tiles (all of which are leaves).
Lemma 3.
A weak tredoku pattern with $\tau$ tiles $\rho$ runs and $\ell$ leaves does not exist when $\ell>\lceil\tau/2\rceil+1$.
This is enough to establish Theorem 2 (i), because a tredoku pattern is a weak tredoku pattern.
Proof.
The statement is true when $\tau=2$, as there are no weak tredoku patterns with $2$ tiles. When $\tau=3$, a weak tredoku pattern has $\rho=1$ and $\ell=3$. Since $\ell\leq\lceil\tau/2\rceil+1$ in this case, the statement is also true when $\tau=3$.
Suppose, for a contradiction, that there exists a weak tredoku pattern with $\tau$ tiles, $\rho$ runs and $\ell$ leaves, with $\ell>\lceil\tau/2\rceil+1$. So $\tau\geq 4$ and, without loss of generality, we may assume that $\tau$ is as small as possible. A run of length $3$ cannot contain $3$ leaves, as the pattern is (edge-)connected and contains at least four tiles. Suppose the pattern has a run containing $2$ leaves. Removing these leaves, we produce a weak tredoku pattern with $\tau^{\prime}$ tiles and $\ell^{\prime}$ leaves, where $\tau^{\prime}=\tau-2$ and (because the non-leaf in our run now becomes a leaf) $\ell^{\prime}=\ell-1$. But $\tau^{\prime}<\tau$, and
$\lceil\tau^{\prime}/2\rceil+1=\lceil\tau/2\rceil>\ell-1=\ell^{\prime}.$
By our choice of $\tau$, we have a contradiction. So all of the $\rho$ runs of length $3$ contain at most one leaf, and thus $\ell\leq\rho$. Since Lemma 1 also holds for weak tredoku patterns, we see that $2\tau=\ell+3\rho\geq 4\ell$ and so $\lceil\tau/2\rceil\geq\ell$. This contradicts the inequality $\ell>\lceil\tau/2\rceil+1$, and the lemma follows. ∎
Note that the proof above does not work if we replace ‘weak tredoku pattern’ by ‘tredoku pattern’ throughout. For removing two leaves in a run can produce a pattern that violates Condition 4 of the tredoku pattern definition: see, for example, the last tredoku pattern in Figure 14.
4 Some consequences of the connectivity conditions
We defined a weak tredoku pattern in Section 3. The aim of this section is to show (Theorem 5) that it is easy to check whether a weak tredoku pattern is in fact a tredoku pattern, by checking the position of leaves within a run. Before this, we show that a weak tredoku pattern does not touch at any corner $p$, in the following sense:
Definition.
Suppose we have pattern $\mathcal{T}$. Let $p$ be a corner of a tile that lies on the boundary of $\mathcal{T}$. The fan of $\mathcal{T}$ at $p$ is the graph $G_{p}$ whose vertices are the tiles in $\mathcal{T}$ with corner $p$, and two tiles are adjacent if they share a common edge. We say that the tiling is singular at $p$ if $G_{p}$ is not connected.
Figure 15 illustrates some of the situations that might arise.
Figure 15: The point $p$ is highlighted, and the tiles from a tiling $\mathcal{T}$ with a corner at $p$ are drawn. In the first three situations, $\mathcal{T}$ is not singular at $p$. In the last three situations, $\mathcal{T}$ is singular at $p$.
Definition.
A pattern $\mathcal{T}$ is nonsingular if for all corners $p$ of tiles in its boundary, $\mathcal{T}$ is not singular at $p$.
Lemma 4.
A weak tredoku pattern $\mathcal{T}$ is nonsingular.
Proof.
Let $p$ be a corner of a tile that lies on the boundary of $\mathcal{T}$. Suppose for a contradiction that $\mathcal{T}$ is singular at $p$. Then the fan $G_{p}$ ot $\mathcal{T}$ at $p$ is not connected. Let $t_{1}$ and $t_{2}$ be tiles that lie in different components of $G_{p}$. Since $t_{1}$ and $t_{2}$ lie in different components, there exist an equilateral triangle $x_{1}$ that is not covered by $\mathcal{T}$, that has $p$ as a corner, and lies between $t_{1}$ and $t_{2}$ as we travel around $p$ clockwise. There exists another such equilateral triangle $x_{2}$ lying between $t_{2}$ and $t_{1}$. See Figure 16 for an illustration of this situation, where the regions $x_{1}$ and $x_{2}$ are shaded. Let $\Gamma(\mathcal{T})$ be the graph whose vertices are the tiles of $\mathcal{T}$, with vertices being adjacent if and only if the tiles share an edge. Since $\mathcal{T}$ is a weak tredoku pattern, $\Gamma(\mathcal{T})$ is connected and so there is a simple path from $t_{1}$ to $t_{2}$ in $\Gamma(\mathcal{T})$. This path induces a path in $|\mathcal{T}|$ from the mid-point of $t_{1}$ to the mid-point of $t_{2}$, by connecting the mid-points of adjacent tiles in the path in $\Gamma$ by line segments. We may complete this path to a non-self-intersecting loop $\pi$ passing through $p$ (the dashed cycle in the figure) by connecting the mid-point of $t_{2}$ to $p$ and then to the mid-point of $t_{1}$. But one of the regions $x_{1}$ and $x_{2}$ lies inside $\pi$, and so $\pi$ cannot be contracted to a point. This contradicts the fact that $|\mathcal{T}|$ is simply connected, as required. ∎
$\scriptstyle t_{1}$$\scriptstyle t_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$
Figure 16: A weak tredoku pattern does not touch at any corner.
Theorem 5.
A weak tredoku pattern is a tredoku pattern if and only if no leaf lies at the centre of its run of length $3$.
The theorem states that we may replace Condition 4 in the definition of a tredoku pattern by the condition that no leaf lies in the centre of its run of length $3$.
Proof.
Let $\mathcal{T}$ be a weak tredoku pattern, and suppose that $t$ is a leaf in $\mathcal{T}$ that lies in the centre of its run $e_{1},t,e_{2}$ of length $3$. We show that $\mathcal{T}$ is not a tredoku pattern.
Let $\Gamma(\mathcal{T})$ be the graph whose vertices are identified with tiles in $\mathcal{T}$, with vertices adjacent when tiles share an edge. As $\mathcal{T}$ is a weak tredoku pattern, $\Gamma(\mathcal{T})$ is connected. The vertex $t$ has degree $2$ in the graph, with adjacent vertices $e_{1}$ and $e_{2}$. We claim that the removal of $t$ disconnects $\Gamma(\mathcal{T})$, because $e_{1}$ and $e_{2}$ lie in separate components. To see this, suppose for a contradiction that there is a path in $\Gamma(\mathcal{T})$ from $e_{1}$ to $e_{2}$ avoiding $t$. By adding $t$ to this path, we produce a cycle in the graph. By connecting mid-points of tiles in the cycle, we find a loop $\pi$ contained in $|\mathcal{T}|$. The tile $t$ is edge-adjacent to two equilateral regions $x_{1}$ and $x_{2}$ not covered by $\mathcal{T}$, corresponding to the sides of the leaf $t$ that are not adjacent to tiles in $\mathcal{T}$. One of these regions is in the interior of our loop $\pi$. But this contradicts that fact that $|\mathcal{T}|$ is simply connected, as it is a weak treduko pattern. So our claim follows.
Let $\mathcal{T}^{\prime}=\mathcal{T}\setminus\{t\}$. We show that $\mathcal{T}$ is not a tredoku pattern by showing that $|\mathcal{T}^{\prime}|$ is not path connected. We first observe that $\mathcal{T}^{\prime}$ is nonsingular. To see this, we first note that since $\mathcal{T}$ is a weak tredoku pattern, Lemma 4 implies that we only need to check that $p$ is not singular for the corners $p$ of $t$. But then $p$ is adjacent to one of the regions $x_{1}$ and $x_{2}$ not covered by $\mathcal{T}$. Moreover, $p$ is adjacent to $t$ and one of the tiles $e_{1}$ and $e_{2}$. So the fan at $p$ remains connected when $t$ is removed, as $t$ lies at the end of the connected component in the fan of $\mathcal{T}$ at $p$. Hence $\mathcal{T}^{\prime}$ is nonsingular.
Suppose, for a contradiction, that $|\mathcal{T}^{\prime}|$ is path connected. Let $\pi$ be a path in $|\mathcal{T}^{\prime}|$ between the mid-points of the tiles $e_{1}$ and $e_{2}$. By the simplicial approximation theorem, we may assume that this path is piecewise linear. The path $\pi$ might pass through corners of tiles in $\mathcal{T}^{\prime}$, but we may modify it so that $\pi$ avoids all corners as $\mathcal{T}^{\prime}$ is nonsingular. Then $\pi$ induces a path from $e_{1}$ to $e_{2}$ in $\Gamma(\mathcal{T})$ that avoids $t$. This is a contradiction, as we have shown above that $e_{1}$ and $e_{2}$ lie in separate connected components of $\Gamma(\mathcal{T})$ after $t$ is removed. Hence $\mathcal{T}^{\prime}$ is not path connected and hence $\mathcal{T}$ is not a tredoku pattern, as required.
We now prove the converse of the theorem. Suppose that $\mathcal{T}$ is a weak tredoku pattern with no leaves in the centre of their runs. We show that $\mathcal{T}$ remains path connected after the removal of any tile $t$. This is sufficient to prove the theorem.
First suppose that $t$ is a leaf. Since $t$ lies at an end of its run, $t$ has degree $1$ in the graph $\Gamma(\mathcal{T})$, and so its removal does not disconnect $\Gamma(\mathcal{T})$. So the pattern $\mathcal{T}^{\prime}$ produced from $\mathcal{T}$ after removing $t$ is still path connected; indeed, $\mathcal{T}^{\prime}$ is even edge connected.
Now suppose that $t$ is not a leaf. Let $\mathcal{T}^{\prime}$ be the pattern produced from $\mathcal{T}$ after removing $t$. Let $\Delta$ be the set of points in the plane shared by $|t|$ and $|\mathcal{T}^{\prime}|$. Clearly, every point $s$ in $|\mathcal{T}^{\prime}|$ is connected by a path to some point in $\Delta$. (We may, for example, truncate a piecewise-linear path in $\mathcal{T}$ from $s$ to the centre of $t$ when it first reaches the boundary of $t$.) To show that $\mathcal{T}^{\prime}$ is path-connected, we just need to show that $\Delta$ is path-connected. If there are four or three tiles in $\mathcal{T}^{\prime}$ adjacent to $t$, the region $\Delta$ is the entire boundary of $t$ or three of the four edges of $t$ respectively, which is a path connected region. If there are just two tiles adjacent to $t$, these tiles are not adjacent to opposite edges of $t$, as $t$ is not a leaf. So $\Delta$ consists of two edges of $t$ that share a common end-point, and again $\mathcal{T}^{\prime}$ is edge connected. This proves the theorem. ∎
Corollary 6.
The leaves in a tredoku pattern lie at the ends of their runs of length $3$.
In the following section, we consider subsets of tiles in a tredoku pattern, and wish to show the region they define is simply connected. The following (standard) lemma provides a simple criterion to check.
Definition.
Let $\mathcal{U}$ be a pattern, and let $\mathcal{U}^{\prime}$ be a subset of the tiles in $\mathcal{U}$. Define $\mathcal{C}$ to be the pattern consisting of all tiles in $\mathcal{U}$ that are not in $\mathcal{U}^{\prime}$. We say that $\mathcal{C}$ is the complement to $\mathcal{U}^{\prime}$ in $\mathcal{U}$.
Lemma 7.
Let $\mathcal{U}$ be a simply connected pattern. Let $\mathcal{U}^{\prime}$ be a subset of the tiles in $\mathcal{U}$. Let $\mathcal{C}$ be the complement to $\mathcal{U}^{\prime}$ in $\mathcal{U}$. Let $\Delta=|\mathcal{U}^{\prime}|\cap|\mathcal{C}|$ be the intersection of the regions defined by $\mathcal{U}^{\prime}$ and $\mathcal{C}$. If $\Delta$ is a single edge, then $\mathcal{U}^{\prime}$ is simply connected.
Proof.
Clearly $\mathcal{U}^{\prime}$ is path connected: paths from any points $x,y\in\mathcal{U}^{\prime}$ to $\mathcal{B}$ can be obtained by truncating paths to a point $\mathcal{C}$, and $\mathcal{B}$ is path connected. Let $\pi$ be a loop in $\mathcal{U}^{\prime}$. By the simplicial approximation theorem, by replacing $\pi$ by a homotopic loop if necessary, we may assume that $\pi$ is piecewise linear. We may contract $\pi$ to a point via some piecewise linear homotopy $h:[0,1]\times[0,1]\rightarrow\mathcal{U}$, since $\mathcal{U}$ is simply connected. The set $\{(x,t)\in[0,1]\times[0,1]:h(x,t)\not\in\mathcal{U}^{\prime}\}$ is a finite set of open polygons, bounded by elements in $\mathcal{B}$. Since $\mathcal{B}$ is a single edge, we may modify the homotopy on each such polygon using a piecewise-linear function in $\mathcal{B}$ to produce a homotopy $h^{\prime}:[0,1]\times[0,1]\rightarrow\mathcal{U}^{\prime}$ in $\mathcal{U}^{\prime}$ that contracts $\pi$ to a point. So $\mathcal{U}^{\prime}$ is simply connected, as required. ∎
5 Classifying tredoku patterns with $\tau=2\rho+1$
Definition.
A tredoku pattern is verdant if it has $12\rho+$ 1 tiles, $\rho+2$ leaves and $\rho$ runs of length $3$.
Verdant tredoku patterns are exactly those whose parameters lie on the uppermost diagonal in Figure 2:
Lemma 8.
Let $\mathcal{T}$ be a tredoku pattern with $\tau$ tiles, $\ell$ leaves and $\rho$ runs of length $3$. Then $\mathcal{T}$ is verdant if and only if $\ell=\lceil\tau/2\rceil+1$.
Proof.
If $\mathcal{T}$ is verdant, it is clear that $\ell=\lceil\tau/2\rceil+1$. Conversely suppose that $\ell=\lceil\tau/2\rceil+1$. Suppose for a contradiction that $\tau$ is even. By Lemma 1,
$\rho=(2\tau-\ell)/3=(2\tau-\tau/2-1)/3=\tau/2-1/3.$
This is a contradiction, since $\rho$ is an integer, and so we may assume that $\tau$ is odd. By Lemma 1,
$\rho=(2\tau-\ell)/3=(2\tau-\tau/2-3/2)/3=(\tau-1)/2$
and so $\tau=2\rho+1$. Thus $\ell=\lceil\tau/2\rceil+1=\rho+2$, and hence $\mathcal{T}$ is verdant, as required. ∎
So verdant tredoku patterns have the largest possible number of leaves given their number of tiles. This section aims (Theorem 11 below) to classify all verdant tredoku patterns. We begin with two methods for building larger tredoku patterns from smaller ones.
Definition.
Let $\mathcal{T}$ be a tredoku pattern, and let $t$ be a leaf in this pattern. Suppose there exist tiles $s_{1}$ and $s_{2}$, not in $\mathcal{T}$, such that $s_{1},t,s_{2}$ is a run of length $3$. Moreover, suppose the intersection of each tile $s_{i}$ with the region $\mathcal{T}$ is contained in $t$. Then $\mathcal{T}\cup\{s_{1},s_{2}\}$ is a $2$-leaf extension of $\mathcal{T}$ at $t$.
Note that many leaves $t$ will not possess tiles $s_{1}$ and $s_{2}$ satisfying the conditions above. For example, in Figure 11 only the shaded leaf possesses neighbouring tiles $s_{1}$ and $s_{2}$ of the right form. In this case, there is a unique choice for the tile northwest of the shaded leaf, but either choice for the tile southeast of the leaf will work. The $2$-leaf extension in this example is also a tredoku pattern, with $7$ tiles and $5$ leaves. As an aside, we state the following lemma, which shows that the $2$-leaf extension of a tredoku pattern is always a tredoku pattern.
Lemma 9.
Let $\mathcal{T}$, $t$, $s_{1}$ and $s_{2}$ satisfy the conditions in the definition above, and let $\mathcal{T}^{\prime}=\mathcal{T}\cup\{s_{1},s_{2}\}$ be the $2$-leaf extension of $\mathcal{T}$. Then $\mathcal{T}^{\prime}$ is a tredoku pattern. If $\mathcal{T}$ has $\tau$ tiles, $\rho$ runs of length $3$ and $\ell$ leaves, then $\mathcal{T}^{\prime}$ has $\tau+2$ tiles, $\rho+1$ runs of length $3$ and $\ell+1$ leaves.
We remark that if $\mathcal{T}$ is verdant, then so is its $2$-leaf extension.
Definition.
Let $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ be tredoku patterns. Suppose that the intersection of the regions $|\mathcal{T}_{1}|$ and $|\mathcal{T}_{2}|$ is exactly $|t|$, where $t$ is a tile which is common to both patterns. Moreover, suppose that $t$ is a leaf in both $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$, the runs in $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are in different directions, and $t$ is at the end of its run in both $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. We say that the pattern $\mathcal{T}=\mathcal{T}_{1}\cup\mathcal{T}_{2}$ is the merging of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ at $t$.
As examples, we see that the last four patterns in Figure 14 are all mergings: the common tile $t$ is taken to be a non-leaf that is the end of both its runs of length $3$. As an aside, we note that the result of merging $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ at $t$ is always a tredoku pattern:
Lemma 10.
Let $\mathcal{T}_{1}$, $\mathcal{T}_{2}$ and $t$ satisfy the conditions in the definition above. The merging $\mathcal{T}$ of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ at $t$ is a tredoku pattern. If $\mathcal{T}_{i}$ has $\tau_{i}$ tiles, $\rho_{i}$ runs of length $3$ and $\ell_{i}$ leaves, then $\mathcal{T}$ has $\tau_{1}+\tau_{2}-1$ tiles, $\rho_{1}+\rho_{2}$ runs of length $3$ and $\ell_{1}+\ell_{2}-2$ leaves.
We remark that if $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are verdant, then so is their merging $\mathcal{T}$.
We are now in a position to classify all verdant tredoku patterns. Our approach will be to show that a verdant tredoku pattern with more than $5$ tiles is either a $2$-leaf extension of a smaller verdant tredoku pattern or is a merging of two smaller verdant tredoku patterns.
Theorem 11.
Up to symmetry, the patterns with $5$, $7$, $9$, $11$, $13$ and $17$ tiles listed in Figures 17 to 20 are the only verdant tredoku patterns.
Proof.
Let $V$ be the set of patterns listed in Figures 17 to 20. It is easy to check that these patterns are all verdant tredoku patterns. We will show there are no other verdant tredoku patterns (up to symmety).
Certainly there are no verdant tredoku patterns with $3$ tiles. A verdant tredoku pattern with $5$ tiles has two runs and $4$ leaves, so each run has $2$ leaves. These runs must intersect in their centre tiles, by Corollary 6. It is now easy to see that the only $5$ tile verdant tredoku patterns, up to symmetry, are listed in Figure 17. So it is sufficient to show that Figures 18 to 20 list all verdant tredoku patterns with $7$ or more tiles.
We note that $V$ is closed under $2$-leaf extensions. Indeed, there is only one $2$-leaf extension of a pattern in $V$, from a $5$-tile pattern to a $7$-tile pattern, as most leaves cannot be used for $2$-leaf extensions. It is also not hard to check that $V$ is also closed under merging, in the sense that if we merge two patterns that lie in $V$ (up to symmetry), the result is also a pattern in $V$ (up to symmetry). To help with this check, the leaves which are suitable for merging because they could lie at the end of a new run of length $3$ are shaded in the figures. Most of these leaves cannot, in fact, be used to merge two patterns in $V$. This is because patterns would intersect outside the leaf, or runs of length $4$ or more would be created.
We note that all patterns in $V$ with $9$ or more tiles are the result of merging two smaller patterns in $V$. We show below that every verdant tredoku pattern with $7$ or more tiles is either a $2$-leaf extension of a verdant tredoku pattern, or is the merger of two verdant tredoku patterns. This suffices to prove the theorem.
Figure 17: All tredoku patterns with $5$ tiles and $4$ leaves, up to symmetry. Shaded leaves are suitable for merging with another pattern. The darkly shaded leaf is also suitable for $2$-leaf extensions.
Figure 18: The unique tredoku pattern with $7$ tiles and $5$ leaves, and the unique tredoku pattern with $9$ tiles and $6$ leaves, up to symmetry. Shaded leaves are suitable for merging with another pattern. No leaves are suitable for $2$-leaf extensions.
Figure 19: Tredoku patterns with $11$ tiles and $7$ leaves. Shaded leaves are suitable for merging with another pattern. No leaves are suitable for $2$-leaf extensions.
Figure 20: Two tredoku patterns with $13$ tiles and $8$ leaves, and one tredoku pattern with $17$ tiles $10$ leaves. No leaves are suitable for merging or $2$-leaf extensions.
Let $\mathcal{T}$ be a verdant tredoku pattern with $7$ or more tiles. So $\mathcal{T}$ has $\rho$ runs of length $3$, and has $12\rho+$ 1 tiles and $\rho+2$ leaves where $\rho\geq 3$. Suppose, as an inductive hypothesis, that the set $V$ contains all verdant tredoku tilings with fewer than $12\rho+$ 1 tiles.
Suppose two $2$-leaf runs intersect. By Corollary 6, these runs intersect in their central tiles. Each of the $5$ tiles in these two runs are either a leaf, or are contained in both these runs. So these $5$ tiles form an edge-connected component. This is a contradiction, since $\mathcal{T}$ is edge-connected and has more than $5$ tiles. So no two $2$-leaf runs intersect.
Suppose a $2$-leaf run $s_{1},t,s_{2}$ intersects the end tile of another run. By Corollary 6, $s_{1}$ and $s_{2}$ are leaves and so $t$ is the non-leaf that lies at the end of another run. Each tile $s_{i}$ shares an edge with $t$, but (since $s_{i}$ is a leaf) does not share an edge with any other tile in $\mathcal{T}$. Moreover, by Lemma 4 the tiles $s_{i}$ do not share a corner with another tile in $\mathcal{T}$, other than those on the common edge with $t$. Define $\mathcal{T}^{\prime}=\mathcal{T}\setminus\{s_{1},s_{2}\}$. We claim that $\mathcal{T}^{\prime}$ is a verdant tredoku pattern. First, all runs in the pattern $\mathcal{T}^{\prime}$ have length $1$ or $3$ (since a run in $\mathcal{T}$ contains none or both of $s_{1}$ and $s_{2}$). Secondly, since $s_{1}$ and $s_{2}$ are leaves (at the end of their run), $\mathcal{T}^{\prime}$ is edge-connected. The intersection of the regions defined by $\mathcal{T}\setminus\{s_{1}\}$ and $\{s_{1}\}$ is a single edge, which is path connected. Since the region $\mathcal{T}$ is simply connected, Lemma 7 shows that $\mathcal{T}\setminus\{s_{1}\}$ is simply connected. Applying this argument again, we similarly see that $\mathcal{T}\setminus\{s_{1},s_{2}\}$ is simply connected, since $\mathcal{T}\setminus\{s_{1},s_{2}\}$ and $\{s_{2}\}$ intersect in a single edge. So the region $\mathcal{T}^{\prime}$ is simply connected, hence $\mathcal{T}^{\prime}$ is a weak tredoku pattern. The leaves in $\mathcal{T}^{\prime}$ are exactly the tile $t$ together with those leaves in $\mathcal{T}$ that lie in $\mathcal{T}^{\prime}$. So all leaves lie at the end of their runs of length $3$, so $\mathcal{T}^{\prime}$ is a tredoku pattern by Theorem 5. Our claim now follows, since $\mathcal{T}^{\prime}$ has $12\rho-$ 1 tiles, $\rho+1$ leaves and $\rho-1$ runs of length $3$ and so it is verdant. Our inductive hypothesis shows that $\mathcal{T}^{\prime}$ lies in $V$. So $\mathcal{T}^{\prime}$ is the second $5$-tile pattern in Figure 17 and $\mathcal{T}$ is the $7$ tile pattern in Figure 18. In particular, $\mathcal{T}$ lies in $V$ as required.
We may now assume that no $2$-leaf run intersects the end of another run. In particular, if we choose any run $a,b,c$, at most one $2$-leaf run intersects it (and the intersection is at $b$).
For $i\in\{0,1,2\}$, define $\rho_{i}$ to be the number of $i$-leaf runs in $\mathcal{T}$. We have $\rho_{0}+\rho_{1}+\rho_{2}=\rho$, and (counting leaves) we have $\rho_{1}+2\rho_{2}=\rho+2$. So $2+\rho_{0}=\rho_{2}$. Since $\rho_{2}>\rho_{0}$, and each 00-leaf run in $\mathcal{T}$ intersects at most one $2$-leaf run, there exists a $2$-leaf run $a,b,c$ that intersects a $1$-leaf run $d,b,e$. Here $a$ and $c$ are leaves, and without loss of generality $d$ is a leaf.
The tile $e$ lies at the end of the run $d,b,e$, and is not a leaf. Suppose $e$ lies at the end of its second run. Define $\mathcal{T}^{\prime}=\mathcal{T}\setminus\{a,b,c,d\}$. We claim that $\mathcal{T}^{\prime}$ is a verdant tredoku pattern. All runs in $\mathcal{T}^{\prime}$ have length $1$ or $3$, because the same is true for $\mathcal{T}$ and because $d,b,e$ is the only run that intersects $\mathcal{T}^{\prime}$ non-trivially and is not contained in $\mathcal{T}^{\prime}$. The tile $e$ is the only tile in $\mathcal{T}^{\prime}$ that shares an edge with a tile in $\mathcal{T}\setminus\mathcal{T}^{\prime}$. So an edge-connected path between two tiles in $\mathcal{T}^{\prime}$ can be modified to avoid $\{a,b,c,d\}$ (by removing tiles $a$, $b$, $c$ and $d$ in a path, and any duplicate copies of $e$). Since $\mathcal{T}$ is edge-connected, the same is true for $\mathcal{T}^{\prime}$. Since $a$, $c$ and $d$ are leaves, the intersection of the regions $\mathcal{T}^{\prime}$ and $\{a,b,c,d\}$ is the common edge of $b$ and $e$, by Lemma 4. By Lemma 7 we find that $\mathcal{T}^{\prime}$ is simply connected. This shows that $\mathcal{T}^{\prime}$ is a weak tredoku pattern. Since $e$ lies at the end of its run in $\mathcal{T}^{\prime}$, we see that $\mathcal{T}^{\prime}$ is a tredoku pattern by Theorem 5. Finally, we note that $\mathcal{T}^{\prime}$ has $32\rho-$ 3 tiles, $\rho$ leaves and $\rho-2$ runs of length $3$ and so $\mathcal{T}^{\prime}$ is verdant and our claim follows. Clearly $\{a,b,c,d,e\}$ is also a verdant tredoku pattern, and so $\mathcal{T}$ is the merging of $\{a,b,c,d,e\}$ and $\mathcal{T}^{\prime}$ at $e$. Since both $\{a,b,c,d,e\}$ and $\mathcal{T}^{\prime}$ lie in $V$ by our inductive hypothesis, and since $V$ is closed under merging, we find that $\mathcal{T}$ lies in $V$, as required.
So we may assume that $e$ lies in the middle of its second run $f,e,g$. The situation we are now considering, up to symmetry, is depicted in Figure 21, where $a$, $c$ and $d$ are leaves, but where $f$ and $g$ may not be.
$e$$b$$d$$f$$g$$a$$c$
Figure 21: Part of a verdant tredoku pattern. Here $a$, $c$ and $d$ are leaves, but $f$ and $g$ may not be. The two shaded regions are not covered by the tredoku pattern.
If $f$ and $g$ are both leaves, we see that $\{a,b,c,d,e,f,g\}$ is an edge-connected component and so forms the whole pattern. In this case, our pattern is the $7$ tile pattern in Figure 18 and so our pattern lies in $V$, as required.
Suppose $f$ is a leaf, but $g$ is not. The tile $g$ lies at an end of the run $f$, $e$, $g$, and is contained in another run of length $3$. Since $a$ is a leaf, $g$ must lie at the end of this second run. Define $\mathcal{T}^{\prime}=\mathcal{T}\setminus\{a,b,c,d,e,f\}$. The same argument as above shows that $\mathcal{T}^{\prime}$ is a tredoku pattern. Indeed, $\mathcal{T}^{\prime}$ has $52\rho-$ 5 tiles, $\rho-1$ leaves and $\rho-3$ runs of length $3$ and so is verdant. So $\mathcal{T}$ is the merging of $\mathcal{T}^{\prime}$ and the $7$ tile verdant treduko pattern $\{a,b,c,d,e,f,g\}$ at $g$. Our inductive hypothesis implies that $\mathcal{T}^{\prime}$ and $\{a,b,c,d,e,f,g\}$ lie in $V$. Since $V$ is closed under merging, we find that our pattern lies in $V$, as required.
When $g$ is a leaf, but $f$ is not, the argument is the same as the paragraph above but with $\mathcal{T}^{\prime}=\mathcal{T}\setminus\{a,b,c,d,e,g\}$ and the merging being at $g$. So in this case, our pattern lies in $V$ as required.
Finally, consider the case when $f$ and $g$ are not leaves. If we remove $e$ from our tiling, the tiles $\{a,b,c,d\}$ form an edge-connected component. Moreover, the tiles $f$ and $g$ are no longer edge-connected. To see this, note that the shaded regions in Figure 21 are not covered by $\mathcal{T}$, since $a$ is a leaf and since the run $d$, $b$. $e$ has maximal length $3$. An edge connected path from $f$ to $g$ that avoids $e$ can be competed to a cycle through $e$ that has one of these shaded regions in its interior. This contradicts that fact that $\mathcal{T}$ is simply connected. So removing the tile $e$ produces three components: the set of tiles connected to each of $d$, $f$ and $g$. We note that the argument using shaded regions above shows that the set of tiles edge-connected to $f$ and the set of tiles edge-connected to $g$ form regions that are disjoint (not even touching at a corner).
Let $\mathcal{T}^{\prime}$ be the set of tiles in $\mathcal{T}$ that are edge-connected to $f$ once we remove $e$ from the tiling. Define $\mathcal{T}^{\prime\prime}=(\mathcal{T}\setminus\mathcal{T}^{\prime})\cup\{f\}$. We claim that both $\mathcal{T}^{\prime}$ and $\mathcal{T}^{\prime\prime}$ are tredoku patterns. The intersection of the patterns $\mathcal{T}^{\prime}$ and $\mathcal{T}^{\prime\prime}$ is exactly the tile $f$, and both patterns are clearly edge-connected. The intersection of $|\mathcal{T}^{\prime}|$ and $|\mathcal{T}\setminus\mathcal{T}^{\prime}|$ is the edge shared between tiles $f$ and $e$. Since $\mathcal{T}$ is edge connected, Lemma 7 implies that $\mathcal{T}^{\prime}$ is simply connected. The tile $f$ lies at the end of both of its runs, so the regions defined by $\mathcal{T}^{\prime\prime}$ and its complement in $\mathcal{T}$ intersect in the edge shared by $f$ and its adjacent tile not equal to $e$. We deduce that $\mathcal{T}^{\prime\prime}$ is also simply connected, by Lemma 7. The only run of length $3$ involving tiles from $\mathcal{T}^{\prime}$ that is not contained in $\mathcal{T}^{\prime}$ is $f,e,g$, so all runs in $\mathcal{T}^{\prime\prime}$ have length $1$ or $3$. Similarly, the only run of length $3$ in $\mathcal{T}^{\prime\prime}$ that is not contained in $\mathcal{T}^{\prime\prime}$ is the run involving $f$ but not $e$. So all runs in $\mathcal{T}^{\prime\prime}$ have length $1$ or $3$. This shows that both $\mathcal{T}^{\prime}$ and $\mathcal{T}^{\prime\prime}$ are weak tredoku patterns. Theorem 5 now implies that $\mathcal{T}^{\prime}$ and $\mathcal{T}^{\prime\prime}$ are tredoku patterns, as claimed. Hence $\mathcal{T}$ is the merging of the tredoku patterns $\mathcal{T}^{\prime}$ and $\mathcal{T}^{\prime\prime}$ at $f$.
We claim that both $\mathcal{T}^{\prime}$ and $\mathcal{T}^{\prime\prime}$ are verdant. Suppose $\mathcal{T}$ contains $\tau$ tiles and has $\ell$ leaves, that $\mathcal{T}^{\prime}$ contains $\tau^{\prime}$ tiles and has $\ell^{\prime}$ leaves, and $\mathcal{T}^{\prime\prime}$ contains $\tau^{\prime\prime}$ tiles and has $\ell^{\prime\prime}$ leaves. We have $\tau=\tau^{\prime}+\tau^{\prime\prime}-1$ and =+′′$\ell=\ell^{\prime}+\ell^{\prime\prime}-2$. Since $\mathcal{T}$ is verdant, we have $\ell=\lceil\tau/2\rceil+1$ and $\tau$ is odd. Suppose, for a contradiction, that $\mathcal{T}^{\prime}$ is not verdant and so $\ell^{\prime}<\lceil\tau^{\prime}/2\rceil+1$. Then
$\ell^{\prime\prime}$ =′′$=\ell-\ell^{\prime\prime}+2$
$>\lceil\tau/2\rceil-\lceil\tau^{\prime}/2\rceil+2$
$\geq(\tau-\tau^{\prime})/2+2\text{, since $\tau$ is odd,}$,
$=\tau^{\prime\prime}/2+3/2$
$\geq\lceil\tau^{\prime\prime}/2\rceil+1.$
But this contradicts Theorem 2(i) (for the tredoku pattern $\mathcal{T}^{\prime\prime}$). Hence $\mathcal{T}^{\prime}$ is verdant. The pattern $\mathcal{T}^{\prime\prime}$ is verdant by the same argument, so our claim follows.
So we see that $\mathcal{T}$ is the merging of verdant tredoku patterns $\mathcal{T}^{\prime}$ and $\mathcal{T}^{\prime\prime}$. Our inductive hypothesis, together with the fact that $V$ is closed under merging, shows that $\mathcal{T}$ lies in $V$. And so the theorem follows. ∎
Corollary 12.
There is no $\tau$ tile, $\ell$ leaf tredoku pattern with $\rho$ runs of length $3$ when $(\tau,\rho,\ell)=(15,7,9)$ or when $(\tau,\rho,\ell)=(2\rho+1,\rho,\rho+2)$ with $\rho\geq 9$.
6 Some unachievable small parameters
In this section, we compete the proof of Theorem 2 by showing that no tredoku pattern exists with $\tau$ tiles, $\rho$ runs of length $3$, and $\ell$ leaves when
$(\tau,\rho,\ell)\in\{(3,1,3),(3,2,0),(4,2,2),(5,3,1),(6,4,0),(12,8,0)\}.$
We begin the section by defining some terminology. First, by a run in a pattern we mean a run of length $3$. Second, we assign a type to each tile in our patttern as follows.
Definition.
Each tile in our pattern is the union of two equilateral triangles joined by a common edge. Two tiles have the same type when these edges are parallel. In Figure 1 the types of tile are shaded differently: We say that the unshaded tiles are of top type, the darkly shaded tiles are of left type and the lightly shaded tiles are of right type.
These names match the apparent direction of the face corresponding to the tile, when we think as the shaded tiles representing a three-dimensional surface made from cubes, with light coming from above and from the right.
Finally we define the direction of a run. Note that the common edges of adjacent tiles in a run are parallel, running in one of three directions. Since each type of tile has edges of two of the three possible directions, any run contains tiles of at most two types.
Definition.
Let $a,b,c$ be a run.
  • The run has direction $LR$ if each of the tiles $a$, $b$ and $c$ are either of left or of right type, and the common edges between adjacent tiles run north–south. We call such a run an $LR$-run.
  • The run has direction $RT$ if each of the tiles $a$, $b$ and $c$ are either of right or of top type, and the common edges between adjacent tiles run southwest–northeast. We call such a run an $RT$-run.
  • The run has direction $LT$ if each of the tiles $a$, $b$ and $c$ are either of left or of top type, and the common edges between adjacent tiles run northwest–southeast. We call such a run an $LT$-run.
The following lemma is easy to prove, and we use it below without comment:
Lemma 13.
Let $\mathcal{T}$ be a pattern.
  • (i)
    The runs in $\mathcal{T}$ in any fixed direction are pairwise disjoint.
  • (ii)
    Any distinct runs in $\mathcal{T}$ are either disjoint, or intersect in a single tile.
The next lemma shows that various small parameter sets are not associated with a tredoku pattern.
Lemma 14.
When
$(\tau,\rho,\ell)\in\{(3,1,3),(3,2,0),(4,2,2),(5,3,1),(6,4,0)\},$
no tredoku pattern $\mathcal{T}$ exists with $\tau$ tiles, $\rho$ runs of length $3$, and $\ell$ leaves.
Proof.
Suppose $(\tau,\rho,\ell)=(3,1,3)$. In this case, $\mathcal{T}$ must consist of a single run, consisting entirely of leaves. But this pattern can be disconnected by removing its middle tile, and so is not a tredoku pattern. (This pattern is a weak tredoku pattern.) So the lemma follows in this case.
Now suppose $(\tau,\rho,\ell)=(3,2,0)$ or $(\tau,\rho,\ell)=(4,2,2)$. Any two distinct runs of length $3$ in a tredoku pattern can intersect in at most one tile. So when $\rho=2$ we must have $\tau\geq 5$. This contradiction shows that the lemma holds in this case.
Suppose that $(\tau,\rho,\ell)=(5,3,1)$. If there are two runs of length $3$ in $\mathcal{T}$ in the same direction, then (since runs in the same direction cannot share a tile) $\tau\geq 6$ and we have a contradiction. So the pattern has exactly one run in each direction. The pattern contains a single leaf, that lies in a single run of length $3$. Consider one of the two runs $a,b,c$ that do not contain the leaf. It contains tiles that are of one of two types, so two of the tiles in $a,b,c$, say $x$ and $y$, are of the same type. Each of $x$ and $y$ is not a leaf, so lies in the run $a,b,c$ and another run $r_{x}$ and $r_{y}$ respectively. These runs are in the same direction, since $x$ and $y$ have the same type. Moreover, $r_{x}$ and $r_{y}$ cannot be the same run, since the only run intersecting $a,b,c$ in two or more positions $x$ and $y$ is $a,b,c$ itself. But our pattern has exactly one run in each direction, and so we have a contradiction. So the lemma also follows in this case.
Finally, suppose that $(\tau,\rho,\ell)=(6,4,0)$. There are $4$ runs of length $3$, so there are at least two runs $r_{1}$ and $r_{2}$ that have the same direction. These runs are disjoint, and so together form a partition of the $6$ tiles in the pattern. Any other run $r$ shares at least two elements with one of $r_{1}$ or $r_{2}$, and so is equal to one of $r_{1}$ and $r_{2}$. This contradicts the fact that we have $4$ runs. Hence the lemma follows. ∎
We now turn our attention to the final set $(12,8,0)$ of parameters. For the rest of this section, $\mathcal{T}$ will be a tredoku tiling with $12$ tiles, $8$ runs and 00 leaves.
Lemma 15.
The tiling $\mathcal{T}$ above has $6$ tiles of one type, and $3$ tiles of each of the other two types. So, by rotating the pattern if necessary, we may assume that $\mathcal{T}$ contains $6$ top tiles, $3$ left tiles and $3$ right tiles.
Proof.
Let $n_{\ell}$, $n_{r}$ and $n_{t}$ be the number of left, right and top tiles respectively, so $n_{\ell}+n_{r}+n_{t}=12$. Since there are no leaves in the pattern, there are exactly $n_{\ell}+n_{r}$ tiles in $LR$-runs. But these runs are disjoint, and each run contains $3$ tiles. So $n_{\ell}+n_{r}$ is divisible by $3$. Hence $n_{t}=12-(n_{\ell}+n_{r})$ is divisible by 3. By counting runs in the other two directions in a similar way, we deduce that $n_{\ell}$ and $n_{r}$ are also divisible by $3$.
Suppose, for a contradiction, that there are no tiles of one type. Without loss of generality, we may assume that $n_{t}=0$. There are $(n_{\ell}+n_{r})/3=4$ $LR$-runs, so there are $8-4=4$ runs in the other two directions $LT$ and $RT$. Each $LT$-run consists entirely of left tiles (as there are no top tiles). Similarly, each $RT$-run consists entirely of right tiles. So each of the $4$ runs not in the direction $LR$ takes one of the following two forms:
or
One of these runs cannot directly be above the other, as these $6$ tiles would lie in distinct $LR$-runs, and there are fewer than $6\times 3=18$ tiles. Now consider the left-most $LT$- or $RT$-run $r_{1}$. Its $3$ tiles are each at the left end of different $LR$-runs. The central tiles of these $LR$-runs must lie in a single $LT$ or $RT$-run $r_{2}$, since we cannot have two such runs above or below the other. Similarly, the right-hand tiles of these $LR$-runs lie in a single $LT$- or $RT$-run $r_{3}$. Let $\mathcal{T}^{\prime}$ consist of tiles in $r_{1}$, $r_{2}$ and $r_{3}$. Each tile in $\mathcal{T}^{\prime}$ lies in two runs that are entirely contained in $\mathcal{T}^{\prime}$. So $\mathcal{T}^{\prime}$ has no leaves: it is an edge-connected component of $\mathcal{T}$. We have a contradiction, since $\mathcal{T}$ is edge-connected and contains $12$ tiles. So we cannot have 00 tiles of any fixed type.
We have that $n_{\ell}+n_{r}+n_{t}=12$ and each summand is a positive multiple of $3$. So (rotating the pattern if necessary) we may assume that $n_{\ell}=n_{r}=3$ and $n_{t}=6$. This proves the lemma. ∎
Lemma 16.
Assume that the pattern $\mathcal{T}$ above has $6$ top tiles, $3$ left tiles and $3$ right tiles. By reflecting the pattern in a vertical line if necessary, we may assume that all $LT$-runs contain $2$ top tiles and $1$ left tile.
Proof.
There are $9$ tiles of top or left type in $\mathcal{T}$. Since each $LT$-run contains $3$ such tiles, and since $LT$-runs are disjoint we see that there are $3$ $LT$-runs in $\mathcal{T}$. Similarly, there are $3$ $RT$-runs and $2$ $LR$-runs in $\mathcal{T}$.
If there are no $LT$-runs consisting entirely of top tiles, then (since the $6$ top tiles must be distributed amongst the $3$ $LT$-runs) the lemma follows. The same is true (after a reflection in a vertical line) if there are no $RT$-runs consisting entirely of top tiles. So we may assume there is an $LT$-run $a,b,c$ and an $RT$-run $d,e,f$ in $\mathcal{T}$ that consist entirely of top tiles. We will aim to derive a contradiction, which proves the lemma.
The tiles $a$, $b$ and $c$ lie in $3$ distinct $RT$-runs, and so one of these $RT$-runs is equal to $d,e,f$. Hence $a,b,c$ and $d,e,f$ intersect. By reflecting $\mathcal{T}$ in a horizontal or vertical line and relabelling if necessary, there are four cases we need to consider: see Figure 22. We derive a contradiction in each case as follows:
$c$$b$$a$$e$$d$
$c$$b$$a$$e$$d$
$c$$b$$a$$d$$f$
$c$$b$$a$$e$$d$
Figure 22: Two runs of top tiles: four cases
Consider the first case in Figure 22. Suppose there is a tile $t$ northeast of $b$. This must be a top tile, as tiles in $\mathcal{T}$ do not overlap. This means we have determined the positions of all top tiles in $\mathcal{T}$, and $\mathcal{T}$ contains the following pattern:
$c$$b$$a$$e$$d$$t$
The $RT$-run through $d$ cannot contain another top tile, and in particular cannot be extended below $d$. Similarly, $LT$-run through $c$ cannot contain another top tile, and cannot be extended above $c$. So we have the following pattern in $\mathcal{T}$:
$c$$b$$a$$e$$d$$t$
This give a contradiction, as the pattern contains parts of $4$ disjoint $LR$-runs (the unlabelled tiles above) and $\mathcal{T}$ contains only $2$ such runs.
Now assume that no tile exists at $t$, and so the $LT$-run through $b$ extends to the southwest. Note that this run cannot contain any other top tiles, as then the pattern would contain at least four $RT$-runs. Arguing similarly for the $RT$-run through $e$, we see that $\mathcal{T}$ must contain the following pattern:
$c$$b$$a$$e$$d$
But this pattern contains parts of four $LR$-runs, and we again have a contradiction, as required.
Now consider the second case in Figure 22. If there is a (top) tile to the northeast of $c$, there is a left tile below $c$, and there are two right-tiles above $d$. We get a contradiction, as $\mathcal{T}$ only contains two $LR$-runs, and we already have parts of three such runs. So there is no tile to the northeast of $c$. No tiles below $a,b,c$ can be top-tiles (as we would have too many $RT$-runs), so we have two left tiles below $c$, in the $LT$-run containing $c$. But the $RT$-run through $e$ cannot extend below $e$, and must contain a right tile. We have parts of three $LR$-runs, and so we have a contradiction.
Consider the third case in Figure 22. Suppose there is a (top) tile to the northeast of $c$. The $LT$-run through $a$ cannot be extended downwards, and contains no more top tiles. So there are two left tiles in the $LT$-run above $a$. Similarly, there are two right tiles in the $RT$-run below $f$. But this means we have parts of four $LR$-runs in $\mathcal{T}$, and so we have a contradiction. So we may suppose there is no top tile to the northeast of $c$ (and so no tile to the southeast of $d$). The tile $c$ is at the upper end of its $LT$-run, and the tile $d$ is at the lower end of its $RT$-run. At most one of the tiles in these runs, other than $c$ and $d$, is a top tile. So again we have a contradiction, as we have parts of three $LR$-runs in $\mathcal{T}$ when we only have two such runs.
Finally, we consider the fourth case in Figure 22. There are no top tiles below $a,b,c$ or $c,e,d$, since we would have too many $RT$- or $LT$-rows respectively. Suppose there is a top tile to the northeast of $b$ (and the northwest of $e$). The $LT$-run through $a$ and the $RT$-run through $d$ contain $4$ non-top tiles, all in disjoint $LR$-runs. This is a contradiction. Now suppose there is a left tile $s$ to the northeast of $b$ and a right tile $t$ to the northwest of $e$. We cannot extend both $a$ and $d$ upwards, as we would produce an $LR$-run of length $4$ through $s,t$. Without loss of generality, $d$ is not extended upwards, and we have the following diagram:
$c$$b$$a$$e$$d$$s$$t$
We have a contradiction, since we have parts of three $LR$-runs. We have a similar contradiction if just one of $s$ and $t$ lies in our pattern. For suppose, without loss of generality, that $s$ lies in the pattern, but $t$ does not. Then $e$ cannot be extended upwards in an $LT$-run, and so we have the following diagram:
$c$$b$$a$$e$$d$$s$
We have parts of three $LR$-runs, giving us our contradiction. Finally, suppose that neither of $s$ nor $t$ lies in our pattern. Then $b$ is at the upper end of its $LT$-run and $e$ is at the upper end of its $RT$-run, and we have the following diagram:
$c$$b$$a$$e$$d$
We have parts of four $LR$-runs, and so we have a contradiction in this case also. This completes the proof of the lemma. ∎
Lemma 17.
Assume that the pattern $\mathcal{T}$ above has $6$ top tiles, $3$ left tiles and $3$ right tiles. Each $LR$-run in $\mathcal{T}$ contains both left and right tiles.
Proof.
By Lemma 16, we may assume without loss of generality that all $LT$-runs contain $2$ top tiles and $1$ left tile.
Suppose, for a contradiction that $\mathcal{T}$ has an $LR$-run with tiles of just one type. Then one $LR$-run $a,b,c$ contains only left tiles, and the other $LR$-run contains only right tiles. The tiles $a$, $b$ and $c$ are in disjoint $LT$-runs, and the $6$ other tiles in these runs are all the top tiles in $\mathcal{T}$. If $a$, $b$ and $c$ lie in the same place within their $LR$-runs, we get a contradiction because because the $9$ tiles form a pattern with no leaves so we cannot add the second $LR$-run to the pattern (see Figure 23). Otherwise, there are top tiles on both sides of $a,b,c$ that are in $RT$-runs that are not entirely made from top tiles. (See Figure 24 for examples). But the right tiles in these runs cannot form a single $LR$-run, and so we have a contradiction as required.
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
Figure 23: Patterns with no leaves.
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
Figure 24: The shaded tiles require their $RT$-runs to be completed with right tiles. Note that shaded tiles lie on both sides of $a,b,c$.
Theorem 18.
No $12$-tile tredoku pattern with 00 leaves exists.
Proof.
Suppose, for a contradiction, that $\mathcal{T}$ is a tredoku pattern with $12$ tiles and 00 leaves. By Lemma 15, we may assume that $\mathcal{T}$ consists of $6$ top tiles, $3$ left tiles and $3$ right tiles. By Lemma 16, we may assume that every $LT$-run contains exactly one left tile. By Lemma 17, $\mathcal{T}$ has two $LR$-runs, with one run $a,b,c$ containing $2$ left tiles and the other run $d,e,f$ containing $1$ left tile.
There are three possibilities for the run $a,b,c$, depending on the location of the right tile. Each such run is drawn in Figures 25 to 27, together with all possibilities for the $LT$-runs intersecting $a,b,c$. (Note that there are exactly $4$ tiles, all top tiles, in these runs. If a combination of runs is not depicted, it can be ruled out because the pattern contains too many $RT$-runs.) Similarly, there are three possibilities for the run $d,e,f$, and all possibilities for this run and the unique $LT$ run intersecting it are drawn in Figures 28 to 30.
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
Figure 25: When $a$ is a right tile.
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
Figure 26: When $b$ is a right tile.
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
$c$$b$$a$
Figure 27: When $c$ is a right tile.
$f$$e$$d$
$f$$e$$d$
Figure 28: When $d$ is a left tile.
$f$$e$$d$
$f$$e$$d$
$f$$e$$d$
Figure 29: When $e$ is a left tile.
$d$$e$$f$
$d$$e$$f$
Figure 30: When $f$ is a left tile.
Since all $LT$-runs and all $LR$-runs in $\mathcal{T}$ are disjoint, we see that the pattern $\mathcal{T}$ is the disjoint union of a pattern $\mathcal{T}_{1}$ from Figures 25 to 27 and a pattern $\mathcal{T}_{2}$ from Figures 28 to 30. But no such combinations give a tiling with the property we want. Here is one way of seeing this. Note that the partial $RT$-runs in $\mathcal{T}_{1}$ and the partial $RT$-runs in $\mathcal{T}_{2}$ must combine to give three $RT$-runs (of length $3$). Reading left-to-right, the lengths of partial $RT$-runs in $\mathcal{T}_{1}$ lie in the set
$\{(2,3),(1,3,1),(3,2),(2,2,1),(1,2,2)\}$
whereas the lengths of partial $RT$-runs in $\mathcal{T}_{2}$ lie in the set $\{(1,2,1),(2,2)\}$. If the sequence of partial runs in $\mathcal{T}_{2}$ is $(1,2,1)$, this is clearly not possible. Suppose the sequence of partial $RT$-runs in $\mathcal{T}_{2}$ is $(2,2)$. In all $3$ cases, the upper tiles in these partial runs touch at a corner, and the same is true for the lower tiles in these runs. The sequence of partial runs in $\mathcal{T}_{1}$ must be $(1,3,1)$. But in all $6$ cases, the tiles in the partial $RT$-runs of length $1$ in $\mathcal{T}_{1}$ do not touch at a corner, so they cannot be combined with the partial runs in $\mathcal{T}_{2}$ to give two $RT$-runs of length $3$. This gives the contradiction we require, and so the theorem follows. ∎
7 Conclusion
  1. 1.
    We introduced the notion of a weak tredoku pattern in Section 3. What parameters are possible for such patterns? Clearly, all the parameters having tredoku patterns are possible. Also, see Figure 31, there are patterns with $\rho$ runs, $12\rho+$ 1 tiles and $\rho$ leaves for all positive integers $\rho$. The proof that there are no tredoku patterns with $12$ tiles and 00 leaves in fact shows there are no weak tredoku patterns with these parameters. Small parameters are easy to deal with. With the results and remarks from Section 3, we have the following theorem:
    Theorem 19.
    Let $\tau$, $\rho$ and $\ell$ be non-negative integers with $\tau\geq 3$, such that $\ell\leq\tau$ and $\ell=2\tau-3\rho$.
    • (i)
      If $\ell>\lceil\tau/2\rceil+1$, then a $\tau$-tile weak tredoku pattern with $\rho$ runs of length $3$ and $\ell$ leaves does not exist.
    • (ii)
      Suppose
      $(\tau,\rho,\ell)\in\{(3,2,0),(4,2,2),(5,3,1),(6,4,0),(12,8,0)\}.$
      Then a $\tau$-tile weak tredoku pattern with $\rho$ runs of length $3$ and $\ell$ leaves does not exist.
    • (iii)
      If neither of the conditions (i) and (ii) hold, then a $\tau$-tile weak tredoku pattern with $\rho$ runs of length $3$ and $\ell$ leaves exists.
    Figure 31: Weak tredoku patterns with $\rho$ runs, $12\rho+$ 1 tiles and $\rho+2$ leaves, for $\rho=1,2,3,4$.
  2. 2.
    Tredoku patterns were originally motivated by tredoku puzzles. Can every tredoku pattern be used as the basis for a tredoku puzzle? So suppose we divide each diamond in a tredoku pattern into a $3\times 3$ grid of sub-diamonds. Define a grouping of sub-diamonds to be the set of sub-diamonds that are contained in a fixed diamond, or the set $9$ sub-diamonds in a ‘straight line’ which follows a run of length $3$. Is there always way of assigning numbers from $\{1,2,\ldots,9\}$ to each sub-diamond so that each grouping of sub-diamonds is a permutation of $\{1,2,\ldots,9\}$? I conjecture the answer is ‘yes’.
  3. 3.
    How does the number of tredoku patterns grow? More precisely, for a positive integer $m$ we may define $f(m)$ to be the number $f(m)$ of tredoku patterns contained within a hexagon with sides of length $m$ (where our equilateral triangles have a side of length $1$). What can be said about the growth of $f(m)$ as $m\rightarrow\infty$? It is interesting to ask the same question of tredoku patterns where we drop the simply connected condition. Here is one natural definition. We define a generalised tredoku pattern to be a pattern of $3$ or more diamond-shaped tiles that satisfies the following conditions:
    1. (a)
      If there exists a run of length $\ell$ in the pattern, then $\ell\in\{1,3\}$.
    2. (b)
      The pattern is edge-connected: tiles are adjacent if and only if they share an edge.
    3. (c)
      The region in the plane defined by the pattern is nonsingular. (In other words, the pattern does not ‘touch at a point’.)
    4. (d)
      If any tile of the pattern is removed, the pattern remains connected. Here tiles are adjacent if they share an edge or a vertex.
    How does the number of generalised tredoku patterns in a hexagon of side $m$ grow as $m\rightarrow\infty$?
Appendix A Donald Preece’s abstract from Combinatorics, Algebra and More, July 2013
Donald A. Preece
(Queen Mary, University of London, UK / University of Kent, UK.)
Tredoku tilings.
We have a supply of identical tiles, each in the shape of a rhombus with alternate angles 60°and 120°, so each tile has 4 sides and 4 vertices. We use $\tau$ of these tiles ($\tau>4$) to form a tiling $\mathcal{T}$ of part of a plane. In $\mathcal{T}$, the position of any vertex of a tile is a vertex of $\mathcal{T}$, and the position of any side of a tile is a side of $\mathcal{T}$
To be a tredoku tiling, $\mathcal{T}$ must have these properties:
  1. 1.
    $\mathcal{T}$ is connected, with no connection consisting merely of two tiles touch- ing at just one vertex;
  2. 2.
    If two tiles in $\mathcal{T}$ touch, either they do so only at a vertex or they have a side in common in $\mathcal{T}$;
  3. 3.
    $\mathcal{T}$ cannot be disconnected by removing just one tile (but a connection consisting merely of two tiles touching at just one vertex is allowed after the removal);
  4. 4.
    If tiles A and B share a side $p$ of $\mathcal{T}$, then there is a third tile C that shares a side $q$ of either A or B, where $q$ is parallel to $p$. This gives a “run” of three tiles.
  5. 5.
    Runs of more than three tiles are not allowed.
  6. 6.
    $\mathcal{T}$ must not have any holes in it.
(There is no requirement for any part of the tiling to repeat.) If a tredoku tiling has $\rho$ runs of three tiles, then
$\frac{3}{2}\rho\leq\tau\leq 2\rho+1.$
Many existence results and constructions are available for tredoku tilings. Motivation: If each tile in $\mathcal{T}$ is subdivided into a $3\times 3$ array of smaller tiles, we have an overall grid for a tredoku© puzzle such as appears daily in The Times.

### v3

@@
 The history of such tilings in combinatorics goes back as least as far as 1896, where Percy MacMahon [4] conjectured a formula for the number of diamond tilings of a hexagonal region of the plane; he established this formula in 1916 [5].
-In physics, diamond tilings are studied as simple instantiations of the 3-dimensional Ising model; in particular there have been some beautiful results on random tilings motivated by this model [2].
+(MacMahon phrased the problem in terms of plane partitions: two-dimensional arrays of integers which are non-decreasing in both rows and columns.
+These objects can be visualised as piles of cubes, the height of each pile corresponding to the integers in the array.
+The appropriate projection of these cubes to a plane gives a diamond tiling.) In physics, diamond tilings are studied as simple instantiations of the 3-dimensional Ising model; in particular there have been some beautiful results on random tilings motivated by this model [2].
 Diamond tilings, and the related domino tilings, are the basic examples which are generalised by the theory of dimers [3].
@@
 In Appendix A, the abstract from [7] is reproduced, which contains the definition in his own words.) In these talks, he provided various constructions for tredoku patterns, and proposed questions and conjectures concerning which parameters are possible.
-Prof Preece died in January 2014, before he was able to complete his work on tredoku patterns.
+Prof.
+Preece died in January 2014, before he was able to complete his work on tredoku patterns.
 This paper solves (in the affirmative) the conjectures he proposed.
-I am working from the notes I took when attending his talk in July 2013 [7].
-Prof.
-Preece distributed copies of his notes on the topic after his talk in October [8], in order to encourage attendees to finish his work.
-Unfortunately, I did not take a copy myself, and I am unable to locate any remaining copies.
-If anyone reading this paper does have a copy of these notes, please get in touch with me: I would be very interested to see their contents.
 In the remainder of the introduction, I will define tredoku patterns, and the parameters we are considering.
@@
 Theorem 2 is visualised by Figure 2, where we plot $\tau$ against $\ell$, and where the shaded squares are parameters where tredoku patterns exist.
-I believe Donald Preece had a proof of Theorem 2 modulo his conjectures above, so the material in Sections 23 and 4 below is a reconstruction of his work.
+Donald Preece had a proof of Theorem 2 modulo his conjectures above, so the material in Sections 23 and 4 below is a reconstruction of his work [^2] After his talk in October [8], Prof.
+Preece distributed copies of his notes on tredoku tilings in order to encourage attendees to finish his work.
+The original version of this paper was submitted with a plea for anyone with these notes to get in touch, and (remarkably!) Martin Ridout contacted me to say that he had a folder containing Prof.
+Preece’s notes on tredoku patterns, including what is almost certainly his handout.
+(Martin was responsible for sorting through Prof.
+Preece’s notes after he passed away.) My thanks to Martin for passing this folder on to me.
+[9].
 His conjectures are solved in Sections 5 and 6.
+After the first version of this paper was submitted, Martin Ridout sent me details of a computer program he had written which enumerates all small tredoku patterns, which in particular shows that the first of Preece’s conjectures holds and provides an alternative proof of Theorem 18 below.
+He hopes to make this data, and a description of the programming techniques, available in the next few months.
 […]
@@
 Figure 10: Tredoku patterns with $15+3i$ tiles and three leaves, for $i=0,1,2$.
-Figure 11: Tredoku patterns with $5$ tiles and $9$ tiles and four leaves. One leaf is shaded in the $9$ tile pattern.
+Figure 11: Tredoku patterns with $5$ tiles and $8$ tiles and four leaves. One leaf is shaded in the $8$ tile pattern.
 Figure 12: Tredoku patterns with $11+3i$ tiles and four leaves, for $i=0,1,2$, with one leaf shaded.
@@
 Proof.
-Clearly $\mathcal{U}^{\prime}$ is path connected: paths from any points $x,y\in\mathcal{U}^{\prime}$ to $\mathcal{B}$ can be obtained by truncating paths to a point $\mathcal{C}$, and $\mathcal{B}$ is path connected.
+Clearly $\mathcal{U}^{\prime}$ is path connected: paths from any points $x,y\in\mathcal{U}^{\prime}$ to $\Delta$ can be obtained by truncating paths to a point $\mathcal{C}$, and $\Delta$ is path connected.
 Let $\pi$ be a loop in $\mathcal{U}^{\prime}$.
@@
 We may contract $\pi$ to a point via some piecewise linear homotopy $h:[0,1]\times[0,1]\rightarrow\mathcal{U}$, since $\mathcal{U}$ is simply connected.
-The set $\{(x,t)\in[0,1]\times[0,1]:h(x,t)\not\in\mathcal{U}^{\prime}\}$ is a finite set of open polygons, bounded by elements in $\mathcal{B}$.
-Since $\mathcal{B}$ is a single edge, we may modify the homotopy on each such polygon using a piecewise-linear function in $\mathcal{B}$ to produce a homotopy $h^{\prime}:[0,1]\times[0,1]\rightarrow\mathcal{U}^{\prime}$ in $\mathcal{U}^{\prime}$ that contracts $\pi$ to a point.
+The set $\{(x,t)\in[0,1]\times[0,1]:h(x,t)\not\in\mathcal{U}^{\prime}\}$ is a finite set of open polygons, bounded by elements in $\Delta$.
+Since $\Delta$ is a single edge, we may modify the homotopy on each such polygon using a piecewise-linear function in $\Delta$ to produce a homotopy $h^{\prime}:[0,1]\times[0,1]\rightarrow\mathcal{U}^{\prime}$ in $\mathcal{U}^{\prime}$ that contracts $\pi$ to a point.
 So $\mathcal{U}^{\prime}$ is simply connected, as required.
@@
 In this case, there is a unique choice for the tile northwest of the shaded leaf, but either choice for the tile southeast of the leaf will work.
-The $2$-leaf extension in this example is also a tredoku pattern, with $7$ tiles and $5$ leaves.
+The $2$-leaf extension in this example is also a tredoku pattern, with $10$ tiles and $5$ leaves.
 As an aside, we state the following lemma, which shows that the $2$-leaf extension of a tredoku pattern is always a tredoku pattern.